# arXiv:2603.02287v1[quant-ph] 02 Mar 2026

### v1

@@ Abstract
In this paper, we employ the quantum regression theorem, a powerful tool in the study of open quantum systems, to analytically study the correlation functions of an Unruh–DeWitt detector, which is an uniformly accelerated two-level quantum system, absorbing charges from an external classical coherent pulse. The system can thus be viewed as a relativistic quantum battery that interacts with the environment of its perceived particles, namely, the quanta of a massless scalar field. By considering the relativistic battery moving in Rindler spacetime, under Born-Markov approximation, we derive the Gorini-Kossakowski-Sudarshan-Lindblad master equation governing the evolution of the system’s reduced density matrix. Moreover, we perform the Fourier transformation of the Wightman functions and use exponential regularisation to compute the functional forms appearing in the master equation. Next, we derive the evolution equations for the the single-time expectation values of the system’s operators. We not only solve these equations to find out the single time averages, but also employ the quantum regression theorem to determine the two-time correlation functions of first and second order. We analyse them in details to explain the phenomenon of spontaneous emission and show analytically how the acceleration can enhance the associated dissipation. Furthermore, we address a special form of second order correlation function relevant to the context of photon bunching arising in Bose-Einstein statistics. Finally, we derive the spontaneous emission spectrum of the battery detector analytically, which in the long-time limit displays a well-defined Lorentzian line shape in the high frequency regime.
preprint: APS/123-QED
@@ I Introduction
The paper is organized as follows. In section II, we provide a brief review on the concept of quantum regression theorem, figuring out the main key steps in its implementation, and presenting examples of evolution equations for the first and second order correlation functions required for our present study of a weakly coupled Markovian system. In section III, we consider a time independent Hamiltonian form of a two level quantum system and couple it linearly to an external classical time dependent pulse. We then construct the system Hamiltonian of the Unruh-DeWitt battery. We also introduce the Hamiltonian of the bath of massless scalar field and present the interaction Hamiltonian as a function of the proper time in the interaction picture. In section IV, we consider the weak-coupling between the system and the bath and derive the evolution equation of the reduced density matrix of the system in the Schrödinger picture. We apply Rindler mapping, parametrize the coordinates of the uniformly accelerating detector in terms of its proper time, and calculate the Wightman function. We then perform the Fourier transform of the Wightman function and determine its real and imaginary part in the dual space of proper time. We also use exponential regularization to extract a finite contribution from the divergent imaginary part. In section V, we define a set of ladder operators for the system Hamiltonian. Next, we calculate the matrix elements of the master equation and determine the evolution of the single time expectation values of the Pauli operators. From these results, we obtain the evolution equations governing the ladder and number operators and solve them to find their single time averages. We also define the fluctuations of those operators with respect to their steady state and present the evolution of the single time averages of those fluctuations. In next section VI, we directly apply QRT and calculate the corresponding fluctuation correlation functions. From these, we compute the first and second order correlators and analyse them to explain various physical phenomena such as spontaneous emission and Hanbury-Brown-Twiss effect. Next, In next section VII, we obtain the power spectrum of the spontaneous emission and analyse the result analytically. Finally, in section VIII, we conclude our study.
@@ II Quantum Regression Theorem: Basic Overview
$\big<\dot{\bm{\hat{\mathcal{A}}}}\big>\,=\,\bm{M}\,\big<\bm{\hat{\mathcal{A}}}\big$>~; (3)
where $\hat{\bm{\mathcal{A}}}$ is a column vector of operators and $M$ is an evolution matrix.
@@
$\Rightarrow\,$ $\dfrac{d}{d\tau}\Big<\hat{\mathcal{O}}(t)\,\hat{\bm{\mathcal{A}}}(t+\tau)\Big>\,=\,\bm{M}\,\Big<\hat{\mathcal{O}}(t)\,\hat{\bm{\mathcal{A}}}(t+\tau)\Big$>~; (4)
@@
The two time correlation functions generated by multiplying a single operator with $\hat{\bm{\mathcal{A}}}$ is known as first order correlation function. However, this order can be chosen to be higher. For example, by multiplying $\hat{\mathcal{O}}^{\prime}(t)$, another operator to the right of $\hat{\mathcal{O}}(t)\,\hat{\bm{\mathcal{A}}}(t+\tau)$, a two time correlation function of second order can be obtained and it governs
@@
$\Rightarrow\,\dfrac{d}{d\tau}\Big<\hat{\mathcal{O}}(t)\,\hat{\bm{\mathcal{A}}}(t+\tau)\,$ $\hat{\mathcal{O}}^{\prime}(t)\Big>\,=\,\bm{M}\,\Big<\hat{\mathcal{O}}(t)\,\hat{\bm{\mathcal{A}}}(t+\tau)\,\hat{\mathcal{O}}^{\prime}(t)\Big$>~. (5)
@@ IV.1 Calculations associated with the Wightman function
$=\frac{\Omega}{2\pi}\coth\left(\frac{\pi\,\Omega}{a}\right)$~ (45)
@@
The point is to note that, although the aforementioned functions provided in eqs. (42, 43, 45, 46) are positive only in the domain $\Omega\,>0$, the following ratios associated with those functions always hold positive value irrespective of the value of $\Omega$. Those are found to be
$\dfrac{\mathcal{G}(\Omega)}{\mathcal{G}_{+}(\Omega)}\,=\,\dfrac{1}{1+e^{-2\pi\Omega/a}}\equiv\,P_{1}(\Omega),$
@@
Now, by fixing $\Omega$ as a positive and finite quantity, we intend to analyse Eqs. (44, 45, 47) with respect to $a$, the acceleration of the battery. While, as per Eq. (44), $\mathcal{G}(0)$ is linearly proportional to the acceleration, $\mathcal{G}{+}(\Omega)$ in Eq. (45) and the associated ratios in Eq. (47) vary exponentially with respect to $a$. By observing the simple exponential structures of these equations, it can be readily recognized that both $\mathcal{G}{+}(\Omega)$ and $\mathcal{G}(-\Omega)\big/\mathcal{G}{+}(\Omega)$ increase with increasing values of $a$. In contrast, $\mathcal{G}(\Omega)\big/\mathcal{G}{+}(\Omega)$ decreases as $a$ increases. This behaviour can be easily verified by substituting numerical values into the relevant equations, which will be very helpful in analysing the nature of the correlation functions in the later sections.
@@ V.2 Matrix elements of GKSL master equation
After considering $\mathcal{K}_{-}(\Omega)\,=\,i\,K_{-}(\Omega)$, the following matrix elements of the master equation easily emerge after substituting eqs. (61, 62, 63) in the above equation. Hence, we have
@@ V.4 Expectation value : Ladder operators and number operator
Taking into account eqs. (64, 65) at first, both the set of equations given in eqs. (80, 81) and in eqs. (82, 83, 84), finally yield
@@
$\dot{\expectationvalue{\bm{\mathcal{S}}}}\,=\,\bm{M}\,\expectationvalue{\bm{\mathcal{S}}}\,+\,\bm{\mathcal{C}}\,;$ (101)
@@
$\bm{\mathcal{S}}\,\equiv\,\begin{pmatrix}\sigma_{-}^{\,(B)}\\ \\ \sigma_{+}^{\,(B)}\\ \\ \sigma_{+}^{(B)}\sigma_{-}^{(B)}\end{pmatrix}\,,\,\bm{\mathcal{C}}\,\equiv\,\begin{pmatrix}0\\ \\ 0\\ \\ \dfrac{\mu^{2}}{4}\mathcal{G}(-\Omega)\end{pmatrix}\,$; (102)
and $\bm{M}\,\equiv$ diag $\Big(M_{11},~M_{22},~\,M_{33}\,\Big)$.
@@
Here we intend to mention a subtle and notable observation regarding this decay. As evident from eqs. (45, 50), the rate of decay $\mathcal{G}_{+}(\Omega)$ depends on $a$, the acceleration of the battery. For a fixed value of $\Omega$, if we regulate the value of uniform acceleration, eq. (45) tells us that the rate of decay increases with the increasing value of $a$ and in the limit $a=0$, $\mathcal{G}_{+}(\Omega)$ takes the lowest value (provided in eq. (50)) for a fixed value of $\Omega$. Hence, dissipation is proportional to the value of the uniform acceleration.
@@
$\dfrac{d}{d\tau}\expectationvalue{\bm{\mathcal{S}}\,+\,\bm{M}^{-1}\,\bm{\mathcal{C}}}\,=\,\bm{M}\,\expectationvalue{\bm{\mathcal{S}}\,+\,\bm{M}^{-1}\,\bm{\mathcal{C}}}$ (108)
@@
$\delta\,\bm{\mathcal{S}}\,\equiv\,\bm{\mathcal{S}}\,+\,\bm{M}^{-1}\,\bm{\mathcal{C}}~.$ (109)
Since $\bm{M}^{-1}\,\bm{\mathcal{C}}$ is a time-independent quantity, therefore, differentiating eq. (109) gives $\dot{\big<\delta\,\bm{\mathcal{S}}\big>}\,=\,\dot{\big<\bm{\mathcal{S}}\big>}$. As also explained in [66], it is actually the expectation value of the operator $\bm{\mathcal{S}}$ when it reaches the steady state and $\dot{\big<\bm{\mathcal{S}}\big>}_{ss}=\,0$. It yields
$\big<\bm{\mathcal{S}}\big>_{ss}\,=\,-\bm{M}^{-1}\,\bm{\mathcal{C}}~;$ (110)
@@
$\delta\,\bm{\mathcal{S}}\,\equiv\,\bm{\mathcal{S}}\,-\,\big<\bm{\mathcal{S}}\big>_{ss}\,=\,\bm{\mathcal{S}}\,+\,\bm{M}^{-1}\,\bm{\mathcal{C}}~.$ (111)
@@
$\dot{\expectationvalue{\delta\bm{\mathcal{S}}}}\,=\,\bm{M}\,\expectationvalue{\delta\bm{\mathcal{S}}}\,;$ (114)
@@
where the matrix elements of $\bm{M}$ are already provided in eqs. (96-98).
@@
In next section, we shall compute two-time correlation functions from eq. (114 or 115) by directly imposing the quantum regression theorem on it.
@@ VI.1 First-order correlation functions
$\big<\dot{\delta\bm{\mathcal{S}}}\big>\,=\,\bm{M}\,\big<\delta\bm{\mathcal{S}}\big$>~, (119)
then according to the quantum regression theorem, the equation of motion for the probability that the operation $\delta\,\bm{\mathcal{S}}(\tau^{\prime}+\tau)$ performed at $(\tau^{\prime}+\tau)$, is affected by an earlier operation $\bm{\mathcal{O}}(\tau^{\prime})$ performed at $\tau^{\prime}$, is given by ($\tau\,\geq\,0$)
$\dfrac{d}{d\tau}\bigg<\mathcal{O}(\tau^{\prime})\,\delta\,\bm{\mathcal{S}}(\tau+\tau^{\prime})\bigg>=\bm{M}\,\bigg<\mathcal{O}(\tau^{\prime})\,\delta\,\bm{\mathcal{S}}(\tau+\tau^{\prime})\bigg$>~; (120)
@@
$\dot{\big<\delta\,\sigma_{-}^{(B)}(\tau^{\prime})\,\delta\,\bm{\mathcal{S}}(\tau+\tau^{\prime})\big>}=\bm{M}\,\big<\delta\,\sigma_{-}^{(B)}(\tau^{\prime})\,\delta\,\bm{\mathcal{S}}(\tau+\tau^{\prime})\big$>~; (122)
$\dot{\Big<\delta\,\sigma_{+}^{(B)}(\tau^{\prime})\,\delta\,\bm{\mathcal{S}}(\tau+\tau^{\prime})\big>}=\bm{M}\,\big<\delta\,\sigma_{+}^{(B)}(\tau^{\prime})\,\delta\,\bm{\mathcal{S}}(\tau+\tau^{\prime})\big$>~; (123)
$\dot{\Big<\delta\,\sum(\tau^{\prime})\,\delta\,\bm{\mathcal{S}}(\tau+\tau^{\prime})\Big>}=\bm{M}\,\Big<\delta\,\sum(\tau^{\prime})\,\delta\,\bm{\mathcal{S}}(\tau+\tau^{\prime})\Big$>~; (124)
@@
We first address eq. (122) which yields a non-zero solution for the second component of $\bm{\mathcal{S}}$. Since, $\delta\,\sigma_{\pm}^{\,(B)}\,=\,\sigma_{\pm}^{\,(B)}$, as shown in eq. (112), the equation concerned reads,
@@
Here, once again, we intend to refer to eqs. (45, 50 ) to remember that, for a finite, positive and fixed value of $\Omega$, the rate of decay with respect to $\tau$ (time delay), increases with the increasing value of uniform acceleration $a$.
@@
Here, it should be mentioned that the above non-zero fraction decreases with the raising acceleration of the battery, as also discussed below eq. (50).
@@
In a similar way, we can also treat eq. (123) which yields the non-zero two time average for the first component of $\bm{\mathcal{S}}$. That is given by
@@
Once again, we intend to remember that the above non-zero fraction increases with the raising acceleration of the battery, as also discussed below eq. (50).
@@
Finally, we attempt to find the last non-zero correlation function of first order from eq. (128) which offers a non-zero solution for the third component of $\bm{\mathcal{S}}$. Therefore,
@@
Nevertheless, the other possible correlations associated with eq. (128) correspond to the first and second components of $\bm{\mathcal{S}}$. Those are obtained as

### v2

# Quantum regression theorem in the Unruh–DeWitt battery

@@ Abstract
In this paper, we employ the quantum regression theorem, a powerful tool in the study of open quantum systems, to analytically study the correlation functions of an Unruh–DeWitt detector, which is an uniformly accelerated two-level quantum system, absorbing charges from an external classical coherent pulse. The system can thus be viewed as a relativistic quantum battery that interacts with the environment of its perceived particles, namely, the quanta of a massless scalar field. By considering the relativistic battery moving in Rindler spacetime, under Born-Markov approximation, we derive the Gorini-Kossakowski-Sudarshan-Lindblad master equation governing the evolution of the system’s reduced density matrix. Moreover, we perform the Fourier transformation of the Wightman functions and use exponential regularisation to compute the functional forms appearing in the master equation. Next, we derive the evolution equations for the the single-time expectation values of the system’s operators. We not only solve these equations to find out the single time averages, but also employ the quantum regression theorem to determine the two-time correlation functions of first and second order. We analyse them in details to explain the phenomenon of spontaneous emission and show analytically how the acceleration can enhance the associated dissipation. Furthermore, we address a special form of second order correlation function relevant to the context of photon bunching arising in Bose-Einstein statistics. We analysed the results both analytically and graphically. Finally, we derive the spontaneous emission spectrum of the battery detector analytically, which in the long-time limit displays a well-defined Lorentzian line shape in the high frequency regime.
@@ I Introduction
The work carried out in this paper are in natural units ($\hbar=1,c=1$) and it is organized as follows. In section II, we provide a brief review on the concept of quantum regression theorem, figuring out the main key steps in its implementation, and presenting examples of evolution equations for the first and second order correlation functions required for our present study of a weakly coupled Markovian system. In section III, we consider a time independent Hamiltonian form of a two level quantum system and couple it linearly to an external classical time dependent pulse. We then construct the system Hamiltonian of the Unruh-DeWitt battery. We also introduce the Hamiltonian of the bath of massless scalar field and present the interaction Hamiltonian as a function of the proper time in the interaction picture. In section IV, we consider the weak-coupling between the system and the bath and derive the evolution equation of the reduced density matrix of the system in the Schrödinger picture. We apply Rindler mapping, parametrize the coordinates of the uniformly accelerating detector in terms of its proper time, and calculate the Wightman function. We then perform the Fourier transform of the Wightman function and determine its real and imaginary part in the dual space of proper time. We also use exponential regularization to extract a finite contribution from the divergent imaginary part. In section V, we define a set of ladder operators for the system Hamiltonian. Next, we calculate the matrix elements of the master equation and determine the evolution of the single time expectation values of the Pauli operators. From these results, we obtain the evolution equations governing the ladder and number operators and solve them to find their single time averages. We also define the fluctuations of those operators with respect to their steady state and present the evolution of the single time averages of those fluctuations. In next section VI, we directly apply QRT and calculate the corresponding fluctuation correlation functions. From these, we compute the first and second order correlators and analyse them to explain various physical phenomena such as spontaneous emission and Hanbury-Brown-Twiss effect. Next, In next section VII, we obtain the power spectrum of the spontaneous emission and analyse the result analytically. Finally, in section VIII, we conclude our study.
@@ II Quantum Regression Theorem: Basic Overview
$\big<\dot{\boldsymbol{\hat{\mathcal{A}}}}\big>\,=\,\boldsymbol{M}\,\big<\boldsymbol{\hat{\mathcal{A}}}\big$>~; (3)
where $\hat{\boldsymbol{\mathcal{A}}}$ is a column vector of operators and $M$ is an evolution matrix.
@@
$\Rightarrow\,$ $\dfrac{d}{d\tau}\Big<\hat{\mathcal{O}}(t)\,\hat{\boldsymbol{\mathcal{A}}}(t+\tau)\Big>\,=\,\boldsymbol{M}\,\Big<\hat{\mathcal{O}}(t)\,\hat{\boldsymbol{\mathcal{A}}}(t+\tau)\Big$>~; (4)
@@
The two time correlation functions generated by multiplying a single operator with $\hat{\boldsymbol{\mathcal{A}}}$ is known as first order correlation function. However, this order can be chosen to be higher. For example, by multiplying $\hat{\mathcal{O}}^{\prime}(t)$, another operator to the right of $\hat{\mathcal{O}}(t)\,\hat{\boldsymbol{\mathcal{A}}}(t+\tau)$, a two time correlation function of second order can be obtained and it governs
@@
$\Rightarrow\,\dfrac{d}{d\tau}\Big<\hat{\mathcal{O}}(t)\,\hat{\boldsymbol{\mathcal{A}}}(t+\tau)\,$ $\hat{\mathcal{O}}^{\prime}(t)\Big>\,=\,\boldsymbol{M}\,\Big<\hat{\mathcal{O}}(t)\,\hat{\boldsymbol{\mathcal{A}}}(t+\tau)\,\hat{\mathcal{O}}^{\prime}(t)\Big$>~. (5)
@@ IV.1 Calculations associated with the Wightman function
$=\frac{\Omega}{2\pi}\coth\left(\frac{\pi\,\Omega}{a}\right)$
$\equiv{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}P_{rod}}$~ (45)
@@
Here we define $\mathcal{G}_{+}(\Omega)$ to be $P_{rod}$ where the suffix “rod” signifies “rate of decay” since $\mathcal{G}_{+}(\Omega)$ often appears as the rate at which dissipation occurs in the single time expectation values and in the correlation functions (as we shall see in section V and VI respectively ) found for the present study. The point is to note that, although the aforementioned functions provided in eqs. (42, 43, 45, 46) are positive only in the domain $\Omega\,>0$, the following ratios associated with those functions always hold positive value irrespective of the value of $\Omega$. Those are found to be
$\dfrac{\mathcal{G}(\Omega)}{\mathcal{G}_{+}(\Omega)}\,=\,\dfrac{1}{1+e^{-2\pi\Omega/a}}\equiv\,P_{1}(\Omega),$ (47)
@@
Now, by fixing $\Omega$ as a positive and finite quantity, we intend to analyse Eqs. (44, 45, 47, 48) with respect to $a$, the acceleration of the battery. As per Eq. (44), $\mathcal{G}(0)$ is linearly proportional to the acceleration. By observing the simple exponential structures of Eq. (47, 48), it can be readily recognized that both $\mathcal{G}{+}(\Omega)$ and $P_{2}(\Omega)$ increase with increasing values of $a$. In contrast, $P_{1}(\Omega)$ decreases as $a$ increases. The behaviour of $P_{1}(\Omega)$, $P_{2}(\Omega)$ ($P_{1}$ and $P_{2}$ respectively) and $P_{rod}$, is graphically represented in Figures. (1, 2, 3), which will be helpful in analysing the nature of the correlation functions in the subsequent section VI.
Figure 1: Behaviour of the decay rate $P_{rod}$ with frequency $\Omega$ for different numerical values of uniform acceleration $a$.
Figure 2: Behaviour of $P_{1}$ with frequency $\Omega$ for different numerical values of uniform acceleration $a$.
Figure 3: Behaviour of $P_{2}$ with frequency $\Omega$ for different numerical values of uniform acceleration $a$.
@@ V.2 Matrix elements of GKSL master equation
After considering $\mathcal{K}_{-}(\Omega)\,=\,i\,K_{-}(\Omega)$, the following matrix elements of the master equation easily emerge after substituting eqs. (62, 63, 64) in the above equation. Hence, we have
@@ V.4 Expectation value : Ladder operators and number operator
Taking into account eqs. (65, 66) at first, both the set of equations given in eqs. (81, 82) and in eqs. (83, 84, 85), finally yield
@@
$\dot{\expectationvalue{\boldsymbol{\mathcal{S}}}}\,=\,\boldsymbol{M}\,\expectationvalue{\boldsymbol{\mathcal{S}}}\,+\,\boldsymbol{\mathcal{C}}\,;$ (102)
@@
$\boldsymbol{\mathcal{S}}\,\equiv\,\begin{pmatrix}\sigma_{-}^{\,(B)}\\ \\ \sigma_{+}^{\,(B)}\\ \\ \sigma_{+}^{(B)}\sigma_{-}^{(B)}\end{pmatrix}\,,\,\boldsymbol{\mathcal{C}}\,\equiv\,\begin{pmatrix}0\\ \\ 0\\ \\ \dfrac{\mu^{2}}{4}\mathcal{G}(-\Omega)\end{pmatrix}\,$; (103)
and $\boldsymbol{M}\,\equiv$ diag $\Big(M_{11},~M_{22},~\,M_{33}\,\Big)$.
@@
Here we intend to mention a subtle and notable observation regarding this decay. As evident from eqs. (45, 51), the rate of decay $\mathcal{G}_{+}(\Omega)$ depends on $a$, the acceleration of the battery. For a fixed value of $\Omega$, if we regulate the value of uniform acceleration, eq. (45) tells us that the rate of decay increases with the increasing value of $a$ and in the limit $a=0$, $\mathcal{G}_{+}(\Omega)$ takes the lowest value (provided in eq. (51)) for a fixed value of $\Omega$. This is also presented in Figure. (1). Hence, dissipation is proportional to the value of the uniform acceleration.
@@
$\dfrac{d}{d\tau}\expectationvalue{\boldsymbol{\mathcal{S}}\,+\,\boldsymbol{M}^{-1}\,\boldsymbol{\mathcal{C}}}\,=\,\boldsymbol{M}\,\expectationvalue{\boldsymbol{\mathcal{S}}\,+\,\boldsymbol{M}^{-1}\,\boldsymbol{\mathcal{C}}}$ (109)
@@
$\delta\,\boldsymbol{\mathcal{S}}\,\equiv\,\boldsymbol{\mathcal{S}}\,+\,\boldsymbol{M}^{-1}\,\boldsymbol{\mathcal{C}}~.$ (110)
Since $\boldsymbol{M}^{-1}\,\boldsymbol{\mathcal{C}}$ is a time-independent quantity, therefore, differentiating eq. (110) gives $\dot{\big<\delta\,\boldsymbol{\mathcal{S}}\big>}\,=\,\dot{\big<\boldsymbol{\mathcal{S}}\big>}$. As also explained in [21], it is actually the expectation value of the operator $\boldsymbol{\mathcal{S}}$ when it reaches the steady state and $\dot{\big<\boldsymbol{\mathcal{S}}\big>}_{ss}=\,0$. It yields
$\big<\boldsymbol{\mathcal{S}}\big>_{ss}\,=\,-\boldsymbol{M}^{-1}\,\boldsymbol{\mathcal{C}}~;$ (111)
@@
$\delta\,\boldsymbol{\mathcal{S}}\,\equiv\,\boldsymbol{\mathcal{S}}\,-\,\big<\boldsymbol{\mathcal{S}}\big>_{ss}\,=\,\boldsymbol{\mathcal{S}}\,+\,\boldsymbol{M}^{-1}\,\boldsymbol{\mathcal{C}}~.$ (112)
@@
$\dot{\expectationvalue{\delta\boldsymbol{\mathcal{S}}}}\,=\,\boldsymbol{M}\,\expectationvalue{\delta\boldsymbol{\mathcal{S}}}\,;$ (115)
@@
where the matrix elements of $\boldsymbol{M}$ are already provided in eqs. (97-99).
@@
In next section, we shall compute two-time correlation functions from eq. (115 or 116) by directly imposing the quantum regression theorem on it.
@@ VI.1 First-order correlation functions
$\big<\dot{\delta\boldsymbol{\mathcal{S}}}\big>\,=\,\boldsymbol{M}\,\big<\delta\boldsymbol{\mathcal{S}}\big$>~, (120)
then according to the quantum regression theorem, the equation of motion for the probability that the operation $\delta\,\boldsymbol{\mathcal{S}}(\tau^{\prime}+\tau)$ performed at $(\tau^{\prime}+\tau)$, is affected by an earlier operation $\boldsymbol{\mathcal{O}}(\tau^{\prime})$ performed at $\tau^{\prime}$, is given by ($\tau\,\geq\,0$)
$\dfrac{d}{d\tau}\bigg<\mathcal{O}(\tau^{\prime})\,\delta\,\boldsymbol{\mathcal{S}}(\tau+\tau^{\prime})\bigg>=\boldsymbol{M}\,\bigg<\mathcal{O}(\tau^{\prime})\,\delta\,\boldsymbol{\mathcal{S}}(\tau+\tau^{\prime})\bigg$>~; (121)
@@
$\dot{\big<\delta\,\sigma_{-}^{(B)}(\tau^{\prime})\,\delta\,\boldsymbol{\mathcal{S}}(\tau+\tau^{\prime})\big>}=\boldsymbol{M}\,\big<\delta\,\sigma_{-}^{(B)}(\tau^{\prime})\,\delta\,\boldsymbol{\mathcal{S}}(\tau+\tau^{\prime})\big$>~; (123)
$\dot{\Big<\delta\,\sigma_{+}^{(B)}(\tau^{\prime})\,\delta\,\boldsymbol{\mathcal{S}}(\tau+\tau^{\prime})\big>}=\boldsymbol{M}\,\big<\delta\,\sigma_{+}^{(B)}(\tau^{\prime})\,\delta\,\boldsymbol{\mathcal{S}}(\tau+\tau^{\prime})\big$>~; (124)
$\dot{\Big<\delta\,\sum(\tau^{\prime})\,\delta\,\boldsymbol{\mathcal{S}}(\tau+\tau^{\prime})\Big>}=\boldsymbol{M}\,\Big<\delta\,\sum(\tau^{\prime})\,\delta\,\boldsymbol{\mathcal{S}}(\tau+\tau^{\prime})\Big$>~; (125)
@@
We first address eq. (123) which yields a non-zero solution for the second component of $\boldsymbol{\mathcal{S}}$. Since, $\delta\,\sigma_{\pm}^{\,(B)}\,=\,\sigma_{\pm}^{\,(B)}$, as shown in eq. (113), the equation concerned reads,
@@
Here, once again, we intend to refer to eqs. (45, 51 ) to remember that, for a finite, positive and fixed value of $\Omega$, the rate of decay with respect to $\tau$ (time delay), increases with the increasing value of uniform acceleration $a$.
@@
Here, it should be mentioned that the above non-zero fraction decreases with the raising acceleration of the battery, as also discussed below eq. (51) and shown in Figure. (2).
@@
In a similar way, we can also treat eq. (124) which yields the non-zero two time average for the first component of $\boldsymbol{\mathcal{S}}$. That is given by
@@
Once again, we intend to remember that the above non-zero fraction increases with the raising acceleration of the battery, as also discussed below eq. (51) and shown in Figure. (3).
@@
Finally, we attempt to find the last non-zero correlation function of first order from eq. (129) which offers a non-zero solution for the third component of $\boldsymbol{\mathcal{S}}$. Therefore,
@@
Nevertheless, the other possible correlations associated with eq. (129) correspond to the first and second components of $\boldsymbol{\mathcal{S}}$. Those are obtained as